### v2

# The effect of predation on the dynamics of Chronic Wasting Disease in deer

The effect of predation on the dynamics of Chronic Wasting Disease in deer
Cody E. FitzGerald1 and James P. Keener2
1 Department of Engineering Sciences and Applied Mathematics, Northwestern University,
Evanston, IL 60208, USA
and
2 Department of Mathematics, University of Utah,
Salt Lake City, Utah, 84112 USA
1 Abstract
Chronic Wasting Disease (CWD) is a neurological disease impacting deer, elk, moose, and other cervid populations and is caused by a misfolded protein known as a prion. CWD is difficult to control due to the persistence of prions in the environment. Prions can remain infectious for more than a decade and have been found in soil as well as other environmental vectors, such as ticks and plants. Here, we provide a bifurcation analysis of a mathematical model of CWD spread in a cervid population, and use a modification of the Gillespie algorithm to explore if wolves can be used as an ecological control strategy to limit the spread of the disease in several relevant scenarios. We then analytically compute the probability that the disease spreads given one infected member enters a fully healthy population and the probability of elimination given a fully susceptible population and remaining prions in the environment. From our analysis, we conclude that wolves can be used as an effective control strategy to limit the spread of CWD in cervid populations, and hunting or other means of lowering the susceptible population are beneficial to controlling the spread of CWD, although it is important to note that inferring biologically relevant parameters from the existing data is an ongoing challenge for this system.
2 Introduction
Chronic Wasting Disease (CWD) is a prion-mediated neurodegenerative disease that has been observed in cervids, such as elk, deer, and moose. Prions are misfolded proteins that can cause neurological disease, such as CWD, but also Creutzfeldt-Jakob disease in humans. CWD is transmitted both directly between susceptible animals and indirectly through environmental exposure. Cervids can become infected through contact with saliva, blood, and other bodily fluids of an infected animal or through indirect environmental infections from water sources, plants, soil, or potentially vector-borne insect infections. CWD is known to influence the movement ecology of deer. A recent study showed that infected deer moved more slowly and were found in lower elevation areas compared with healthy deer that were not starving, and when compared to starving deer, the infected deer more often utilized areas near streams and with lower herbaceous biomass [6]. In recent years, wild populations of cervids have exhibited high levels of CWD prevalence. The 2024 Chronic Wasting Disease Surveillance Report from the Wyoming Game and Fish Department reported that the prevalence of CWD in some populations of mule deer is as high as 66.3%  [^1] The report can be found here https://wgfd.wyo.gov/wyoming-wildlife/wildlife-disease-and-health/chronic-wasting-disease using the drop-down Report menu at the bottom of the website.. CWD has proven difficult to control or eradicate, due to the time scale associated with the persistence of CWD prions in the environment. Prions decay slowly and can remain latent and infectious for more than a decade [12, 39]. Many mathematical models have been developed to help understand the disease spread in deer populations, evaluate the efficacy of control strategies, and infer the modes of transmission. We briefly review several of the published models, but our summary should not be considered exhaustive.
One of the earlier discrete population models examined the spread of CWD and control strategies [25, 15], but the modeling assumptions of the models, in particular frequency-dependent transmission, were reevaluated in the follow-up paper [33]. Miller and co-authors developed and attempted parameter estimation of competing SIR-type models that describe transmission of CWD in a population of cervids through different mechanisms—though at least one of the models considered appears to exhibit a lack of structural identifiability, which we briefly comment on in the Discussion [24]. They found support for indirect transmission of CWD using a model selection approach [24]. A different study also suggests that the disease will spread via indirect transmission due to the persistence of prions in the environment [4]. The model presented in [24] was further modified for further mathematical and control-based investigations.
Sharp and Pastor modified the model presented in [24] to include logistic growth in the susceptible population and performed a mathematical analysis of the system [34]. Wild and co-authors later extended the model presented in [24] to include the effects of predation from wolves [40] and analyzed the system in a non-oscillatory parameter regime, using parameter values estimated in [24] and other studies. An earlier report to the National Park Service by Hobbs also analyzed wolf-predation as a control strategy using a similar deterministic model in an oscillatory parameter regime [16]. Barbera and Pollino performed a mathematical analysis of a PDE [5] that was inspired by the model presented in [34]. Reyes and co-workers designed a spatiotemporal causal inference scheme to examine the effects of culling strategies using a PDE-based approach [22]. Potapov and co-authors used a complex age-structure differential equation model to examine the influence of harvesting strategies on population dynamics and disease suppression [30]. Vasilyva and co-authors performed a mathematical analysis of a differential equation model that describes CWD spread in a population of deer, including both indirect and direct transmission [37]. They also attempted a parameter estimation of the system, using upper and lower bounds for parameter estimation based on estimates of parameters found in [24]. A different study examined the influence of culling strategies to control CWD spread using a differential equation model that incorporated seasonality into a population model of cervids [29]. Others created an integrodifference equation model that incorporated direct and indirect transmission and long-distance dispersal from juvenile deer to understand CWD spread in a population of white-tailed deer [23]. Several groups have examined harvest strategies using mathematical models that delineate the deer population by sex [3, 2, 31]. Control investigations into CWD transmission in deer populations are summarized in [36].
Recently, agent-based models, statistical models, and machine learning approaches have been used to examine CWD transmission in deer populations. An agent-based model has been used to examine the spread of CWD in deer populations and possible management strategies to limit CWD [35]. A recent long-term statistical analysis used movement data collected from 596 deer to examine how seasonal behaviors, habitat selection, and home range size influence CWD transition in southwest Wisconsin [13]. Machine learning approaches have also been applied to the spread of CWD in deer [1].
In this paper, we present a relatively simple model to understand the CWD spread in a population of deer and explore the role of predators, such as wolves, as an ecological control strategy. Our model is inspired by previous mathematical models [34, 40, 16]. We begin by performing a bifurcation analysis to demonstrate the effect of predation on the dynamics of the disease and examine a stochastic model of the same dynamics. We show that the presence of wolves in all cases lessens the severity of a CWD outbreak and can lead to the elimination or at least significant reduction of the severity of the disease in the deer population. This is true even if the control intervention is applied after CWD has spread in the population. The mechanism of this reduction is easy to understand, as predation increases the death rate of infected animals and thereby decreases the amount of spread of prions from the infected animals. We also analytically compute the probability that a fully susceptible population with no infected members will become infected through indirect environmental transmission and show that the probability of infection recurrence scales exponentially with the size of the susceptible population. This analysis shows that CWD transmission in deer may be controlled or limited by a combination of predator-based control strategies, which lower the infected population, and hunting or other harvesting strategies to limit the size of the susceptible population.
3 A Continuous Variable Population Model
The population dynamics model that we begin with was first presented and studied in [34], and is given by
$\frac{dS}{dt}$ $=$ $rS(1-\frac{S}{K})-\gamma SE,$ (1)
$\frac{dI}{dt}$ $=$ $\gamma SE-\mu_{i}I,$ (2)
$\frac{dE}{dt}$ $=$ $\epsilon I-\mu_{e}E.$ (3)
Here, $S$ is the density of susceptible and uninfected animals, $I$ is the density of infected animals, and $E$ is the density of prions in the environment (primarily in the soil). The meaning of the terms of the model is that $S$, in the absence of prions, grows according to logistic dynamics, but becomes infected at a law of mass action rate from prions. The infected animals have a death rate $\mu_{i}$. Prions are shed by infected animals at rate $\epsilon$, and degrade at rate $\mu_{e}$.
A slightly earlier model [40] included the effects of predation using the equations
$\frac{dS}{dt}$ $=$ $r(S+I)(1-\frac{S+I}{K})-S(\gamma E+m)-\delta\rho_{s}(S,I)S,$ (4)
$\frac{dI}{dt}$ $=$ $\gamma SE-(\mu_{i}+m)I-\delta\rho_{i}(S,I)I,$ (5)
$\frac{dE}{dt}$ $=$ $\epsilon I-\mu_{e}E.$ (6)
where $\rho_{s}(S,I)={S+I\over S+vI}$, $\rho_{i}(S,I)=v(1-c){S+I\over S+vI}$.
The analysis of the model in [34] showed that there are three regions of parameter space with distinct dynamic behaviors, a region with no infected animals or prions, a region in which the disease is endemic, but steady, and a region in which there are periodic (oscillatory) outbreaks of the disease, during which the animal population falls catastrophically low. They did not study the effect of predation on the dynamics of the disease. A rigorous mathematical analysis of a closely related model is given in [21]. The model in [40] was used to study the effects of predation through the parameter $\delta$ but failed to notice the existence of periodic solutions and the effect of predation on those oscillations.
The model considered here uses a slightly different density dependence on birth and death and a simplified predation effect, motivated by simple predator-prey interactions,
$\frac{dS}{dt}$ $=$ $bS(1-\frac{S+I}{K^{\prime}})-dS-\gamma_{e}SE-\gamma_{i}SI-\rho_{s}SW,$ (7)
$=$ $rS(1-\frac{S+I}{K})-\gamma_{e}SE-\gamma_{i}SI-\rho_{s}SW,$
$\frac{dI}{dt}$ $=$ $\gamma_{e}SE+\gamma_{i}SI-\mu_{i}I-\rho_{i}IW,$ (8)
$\frac{dE}{dt}$ $=$ $(\epsilon_{i}+\epsilon_{d}\mu_{i}+\epsilon_{w}\rho_{i}W)I-\mu_{e}E\equiv\epsilon(1+\xi W)I\ -\mu_{e}E,$ (9)
where $W$ is the density of the predator (wolf) population, assumed here to be constant, and $\rho_{s}$ and $\rho_{i}$ are both constant, quantifying the effect of predation. Here $b$ is the zero population birthrate, $K^{\prime}$ is the population density at which the birthrate is zero, $d$ is the natural death rate of healthy animals, and r=bdr=b-d, $K=(1-{d\over b})K^{\prime}$, are both assumed to be positive. Further, the model incorporates the assumption that infected animals contribute to the carrying capacity, but not to the birth of susceptible animals. The disease is assumed to spread in one of two ways, through contact with prions in the environment (rate $\gamma_{e}$) and through direct contact between infected and susceptibles (rate $\gamma_{i}$). Prions are deposited into the environment because of shedding by infected animals (at rate $\epsilon_{i}I$) or death of infected animals (at rate $\epsilon_{d}\mu_{i}I$), or possibly from wolves ingesting prions from infected animals and depositing them into the environment through their feces (at rate $\epsilon_{w}\rho_{i}WI$). It is reasonable to assume that $\epsilon_{w}\leq\epsilon_{d}$, since death by predation is surely to produce no more prions than death by disease. Notice also, that the units for $\epsilon_{i}$ (yr-1(prion density)(deer density)-1) are different than those for $\epsilon_{d}$ and $\epsilon_{w}$ ( (prion density)(deer density)-1). We take $\xi W=0$ throughout our analysis. Parameter values used for this study are shown in Table 1, although the value for $\gamma_{i}$ is not well established and is quite uncertain. Without loss of generality, $\gamma_{e}$ can be taken as one by rescaling the units on $E$.
[…]
Table 1: a) birth rate at population =0 (=1.8 fawns per female deer per year) b) death rate = 0.1 (probability of survival per year = 0.9).
It is likely that $\rho_{s}<\rho_{i}$, since wolves prey primarily on aged and disabled deer, and the percentage of such among the susceptible population is certainly smaller than among the CWD-infected deer. In fact, infected deer are noticeably less reactive to threats and thus much easier targets of predation than uninfected deer.
There are three possible steady solutions for this model: the extinct state with $S=0$, the disease-free state, and the disease endemic state. The disease-free state has
$S=K(1-{\rho_{s}\over r}W),\quad I=0,\quad E=0.$ (10)
This solution requires that $\rho_{s}W<r$, i.e., that predation of the wolves on healthy animals is not too large.
The disease endemic state is given by
$S$ $=$ ${\mu_{e}\over\gamma}(\mu_{i}+\rho_{i}W),$ (11)
$I$ $=$ ${\mu_{e}\over\gamma}{r(K\gamma-\mu_{e}\mu_{i})-(K\gamma\rho_{s}+\mu_{e}r\rho_{i})W\over K\gamma+r\mu_{e}},$ (12)
$E$ $=$ ${\epsilon(1+\xi W)\over\mu_{e}}I,$ (13)
where $\gamma=\epsilon\gamma_{e}(1+\xi W)+\gamma_{i}\mu_{e}$. Clearly, this solution only exists provided $r(K\gamma-\mu_{e}\mu_{i})-(K\gamma\rho_{s}+\mu_{e}r\rho_{i})W>0$.
One can also show that the disease-free state is stable if $r(K\gamma-\mu_{e}\mu_{i})-(K\gamma\rho_{s}+\mu_{e}r\rho_{i})W<0$, and unstable if not. It is immediately apparent that the presence of predators decreases the likelihood of an endemic disease state existing and stabilizes the disease-free state. Indeed, if $W$ is large enough, there is only the disease-free state. Furthermore, in the disease-free state, the presence of predators decreases the $S$ population (see (10)), whereas in the endemic disease state, if it exists, the presence of predators increases the $S$ population, while it decreases the level of the infected population $I$ as well as the environmental load of prions, $E$ (see (11)).
Using XPP, Sharp and Pastor [34] showed that there is a curve in parameter space at which there is a Hopf bifurcation for the equations (1-3). It is a direct calculation (i.e., find the location in parameter space at which the Jacobian of the system (1-3) has purely imaginary eigenvalues [18]) to show that this curve is given by
$H(\kappa)=\kappa^{2}-\kappa(\mu_{e}^{2}+3\mu_{e}\mu_{i}+\mu_{i}^{2})-\mu_{e}\mu_{i}r(\mu_{e}+\mu_{i})=0,$ (14)
where $\kappa=K\epsilon\gamma$. Numerical simulations verify that there are stable periodic solutions for parameter values for which $H(\kappa)>0$. Since $H(\kappa)=0$ has a unique positive root, this implies that there are periodic solutions for $\kappa$ sufficiently large.
We can also find the curve of Hopf bifurcation points for equations (7-9). It is given by $H_{W}=0$, where
$H_{W}$ $=$ $K^{3}\epsilon\gamma_{e}(1+\xi W)\gamma^{2}(r-\rho_{s}W)(\gamma-M\gamma_{i})$ (15)
$-K^{2}\gamma r(\gamma_{i}\mu_{e}(2M^{2}\gamma_{i}\mu_{e}-3\gamma M^{2}-3\gamma M\mu_{e})+\beta\gamma^{2})$
$-K\mu_{e}r^{2}(\gamma_{i}\mu_{e}(M^{2}\gamma_{i}\mu_{e}-3\gamma M^{2}-2\gamma M\mu_{e})+\beta\gamma^{2})$
$-Mr^{3}\mu_{e}^{3}(\gamma-M\gamma_{i}),$
where $\beta=M^{2}+(3\mu_{e}+r-\rho_{s}W)M+\mu_{e}^{2}$, $\gamma=\epsilon\gamma_{e}(1+\xi W)+\gamma_{i}\mu_{e}$, $M=\mu_{i}+\rho_{i}W$.
Figure 1: Regions for Chronic Wasting Disease: I: No disease, II: Disease is endemic, III: Disease has oscillatory outbreaks, IV: Extinction (not shown here), where $\frac{\rho_{s}}{\rho_{i}}=0.5$. The curve separating region I from region IV is $\rho_{s}W=r$; The curve separating region I from region II is $r(\kappa-\mu_{i}\mu_{e})-W(\mu_{e}r\rho_{i}+\kappa\rho_{s})=0$, and the curve separating region II from region III is $H_{W}(\kappa)=0$ (15). Asterisks show locations of the solutions shown in Fig. 3 with $K=30$, and $\rho_{i}W=0.0,0.1,0.15,0.5$.
These features of predation are depicted in Fig. 1. There it is shown that parameter space is divided into four regions, I, in which there is no disease, II, in which the disease is endemic, III, in which the disease is oscillatory, and IV, in which the deer population is extinct because of over-predation. The boundaries between these regions are shown plotted in $\rho_{i}W-K$ parameter space, showing the stabilizing effect of predators.
Refer to captionRefer to caption
Figure 2: Bifurcation Diagram for Chronic Wasting Disease, corresponding to a vertical slice of Fig. 1 at $K=30$: I: No disease, II: Disease is endemic, III: Disease has oscillatory outbreaks, IV: Predation leads to extinction. Green curves denote the maximum and minimum of stable oscillations, red curves denote stable steady solutions, and black dashed curves denote unstable steady solutions. For these plots, $\frac{\rho_{s}}{\rho_{i}}=0.5$, $\mu_{e}=0.2$.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 3: Solution of the differential equation system with (Upper Left) $\rho_{i}W=0$, (oscillatory outbreaks), (Upper Right) $\rho_{i}W=0.1$ (suppressed oscillations), $\rho_{i}W=0.15$, and $\rho_{i}W=0.5$ (no oscillations, stable endemic state). For these plots, $\frac{\rho_{s}}{\rho_{i}}=0.5$, $K=30$, $\mu_{e}=0.2$.
Not only does the presence of a predator decrease the likelihood of disease, but it also reduces the prion load of the environment and decreases the amplitude, hence the severity, of oscillatory outbreaks when they occur. This can be seen in numerical simulations, as demonstrated in Figs. 3. Specifically, in these plots are shown the numerical simulation of the ODE system for the fixed value of $K=30$ for different values of $\rho_{i}W=0,0.1,0.15,0.5$, placing these in regions II and III. (See asterisks in Fig. 1). Noticeably, the effect of predation is to reduce the severity of the disease, at the cost of reducing the overall population size below its carrying capacity.
4 Stochastic simulation
Because deer populations are discrete and not too large, a continuous variable model may be inappropriate, and there might be insight to be gained from stochastic simulations of this process. To this end, we consider the following seven reactions:
$R_{1}:$ $s\rightarrow s+1\quad{\rm(Birth~of~deer)}$ (16)
$R_{2}:$ $s\rightarrow s-1\quad{\rm(Death~of~deer)}$ (17)
$R_{3}:$ $s,i\rightarrow s-1,i+1\quad{\rm(Infection~of~deer~by~E)}$ (18)
$R_{4}:$ $s,i\rightarrow s-1,i+1\quad{\rm(Infection~of~deer~by~}i)$ (19)
$R_{5}:$ $i\rightarrow i-1\quad{\rm(Death~of~infected~deer)}$ (20)
$R_{6}:$ $s\rightarrow s-1\quad{\rm(Predation~of~uninfected~deer)}$ (21)
$R_{7}:$ $i\rightarrow i-1\quad{\rm(Predation~of~infected~deer)}$ (22)
where $s$ and $i$ represent the number of susceptibles and infected, respectively, at any given time. (Since we are taking the wolf population to be constant, it is not necessary to include wolf dynamics here.) We need to get the reaction rates correct for the stochastic simulation. We start with some representative area $A$ that we want to consider in our simulation, and realize that the unit of densities is set by some area $A_{0}$, in this case, $A_{0}=100$km2. This means, for example, that if $K$ is the carrying capacity density, then $K$ is also the number size of the carrying capacity in an area of size $A_{0}$. Therefore, the size of the carrying capacity for an area of size $A$ is $K{A\over A_{0}}$. Now, let $s$ and $i$ be the integers that the densities $S$ and $I$ represent. Then, if $S$ is the density of susceptibles in the area $A_{0}$, the number of susceptibles in the area $A$ is $s=S{A\over A_{0}}$, and similarly for infected animals, $i=I{A\over A_{0}}$.
The reaction rates for the stochastic process are, therefore,
$R_{1}:$ $r_{1}=bs(1-{A_{0}\over K^{\prime}A}(s+i)),$ (23)
$R_{2}:$ $r_{2}=ds,$ (24)
$R_{3}:$ $r_{3}=\gamma_{e}Es,$ (25)
$R_{4}:$ $r_{4}=\gamma_{i}{A_{0}\over A}si,$ (26)
$R_{5}:$ $r_{5}=\mu_{i}i,$ (27)
$R_{6}:$ $r_{6}=s\rho_{s}W,$ (28)
$R_{7}:$ $r_{7}=i\rho_{i}W,$ (29)
Here, $E$ remains a continuous, not discrete, variable, and so is governed by the differential equation $\frac{dE}{dt}=\frac{\epsilon A_{0}}{A}i-\mu_{e}E$. We would like to use these reaction rates with the Gillespie algorithm [14] to do stochastic simulations. However, since $E$ is continuously changing in time, the Gillespie algorithm is not directly applicable. Instead, we use “Poisson thinning” to account for this fact, as follows[20, 28, 32]: Since between reactions, $E(t)$ is a monotone function, bounded between $E(t_{0})$ and $\epsilon\frac{iA_{0}}{\mu_{e}A}$, where $t_{0}$ is the last reaction time, we let $E^{*}={\rm max}[E(t_{0}),\epsilon\frac{iA_{0}}{\mu_{e}A}]$, and introduce a modified reaction $R_{3}^{*}$ (since $R_{3}$ depends on $E$), with reaction rate $r_{3}^{*}=\gamma sE^{*}$ (which is the fastest possible reaction rate), and then proceed with the usual Gillespie determination of next reaction time and next reaction. However, whenever the next reaction is to be $R_{3}^{*}$ at time $t^{*}$, the reaction is split into two, only one of which is implemented: a null reaction for which no change to $s$ or $i$ is made with probability $\frac{E^{*}-E(t^{*})}{E^{*}}$, and the reaction that reduces $s$ by one and increases $i$ by one with probability $\frac{E(t^{*})}{E^{*}}$.
Because $K$ is in the range of 0-100(100 km2)-1, so $A_{0}=100$km2, to have a reasonable number of deer, it must be that A=1000A=1000km2, at least. This also needs to take into account the typical range of a wolf pack of between 130 -2600 km2.
Refer to captionRefer to caption
Figure 4: Example of stochastic simulation showing (Left) $s$ (blue) and $i$ (red) as a function of time, compared with the deterministic solution, shown dashed, starting with one infected cervid and no prions at time $t=0$, and (Right) $E(t)$ vs. $t$ for 50 different trials. Parameter values are K=30K=30(100km2)-1, $A=2000$ km2, $\mu_{e}=0.2$yr-1, $\rho_{i}W=0$.
Refer to captionRefer to caption
Figure 5: Example of stochastic simulation showing (Left) $s$ (blue) and $i$ (red) as functions of time, compared with the deterministic solution, shown dashed, starting with one infected cervid and no prions at time $t=0$, and (Right) $E(t)$ vs. $t$ for 50 different trials. Parameter values are K=30K=30(100km2)-1, $A=2000$ km2, $\mu_{e}=0.2$yr-1, $\rho_{i}W=0.1$.
Refer to captionRefer to caption
Figure 6: Example of stochastic simulation showing (Left) $s$ (blue) and $i$ (red) as functions of time, compared with the deterministic solution, shown dashed, starting with one infected cervid and no prions at time $t=0$, and (Right) $E(t)$ vs. $t$ for 50 different trials. Parameter values are K=30K=30(100km2)-1, $A=2000$ km2, $\mu_{e}=0.2$yr-1, $\rho_{i}W=0.5$.
Refer to captionRefer to caption
Figure 7: Example of stochastic simulation showing (Left) $s$ (blue) and $i$ (red) as functions of time, compared with the deterministic solution, shown dashed, starting with one infected cervid and no prions at time $t=0$, and (Right) $E(t)$ vs. $t$ for 50 different trials. Parameter values are K=30K=30(100km2)-1, $A=2000$ km2, $\mu_{e}=0.2$yr-1, $\rho_{i}W=0.75$.
The first stochastic simulation shown (Fig. 4) is with no predators $\rho_{i}W=0$, and K=30K=30(100km2)-1, and $\mu_{e}=0.2$yr-1. For these parameter values, the deterministic model has oscillatory solutions. For this model with the stochastic simulation, the population $s$ always goes extinct in finite time [8], and so to prevent that from happening, $s$ is artificially never allowed to become zero. In Fig. 4(Right) are shown the prion density trajectories for 50 independent trials, and it is seen here that most of the invasions with one infected cervid result in an oscillatory prion epidemic. The Figs. 5-7 show a sample trajectory (Left) and the prion density for 50 trajectories (Right) for the three values of $\rho_{i}W=0.1,0.5$, and 0.75.
There are several observations to make from these simulations. First, the probability of an epidemic getting started from one infected individual is a decreasing function of the number of wolves present at the time of the appearance of the infected individual. Second, it could be that the diseased animals are eradicated, but the disease reemerges because of the remnant prions in the environment. However, third, there is a possibility that the disease is permanently eradicated even after it initially spreads, and this probability is an increasing function of the number of wolves present. For example, in Fig. 6 roughly 80% of the trajectories show disease elimination (with $\rho_{i}W=0.5$), while with $\rho_{i}W=0.75$ (Fig. 7) about 98% of the trajectories show elimination of the disease.
More data for the probability of disease eradication is shown in Fig. 8, where the probability of long-term disease survival as a function $\rho_{i}W$ is shown, determined using 200 trajectories for each value of $\rho_{i}W$.
Figure 8: Probability of long-term survival of the disease following infection of a single individual, plotted as a function of $\rho_{i}W$.
4.1 Disease Control
For the previous simulations, it was assumed that there was a preexisting population of predators at the time the infected animal was introduced. However, for a variety of reasons, this may not be the case in a real situation. Instead, it may be that wolves can be introduced only after the disease has been introduced and has begun to spread. Thus, at the time of discovery, wildlife managers may introduce a population of predators and wonder if this is able to bring the disease under control. For this simulation, we assume that the disease is discovered when there are $\frac{1}{20}$th of the carrying capacity infected animals, at which time wolves are introduced. The result without introduction of predators is that shown in Fig. 4. The result of this introduction is summarized in Figs. 9-10. The phase portrait course of the disease for 50 trajectories with no intervention is shown in Fig. 9(Left), and with the introduction of wolves with $\rho_{i}W=0.8$ in Fig. 9(Right) and with $\rho_{i}W=1.5$ in Fig. 10(Left). In Fig. 10(Right) is shown the probability of disease survival as a function of time for these three scenarios. Apparent from these plots is that with the introduction of wolves, the severity of the disease is decreased in two ways, namely, the maximal level of infected animals is decreased and the survival of uninfected animals is increased, giving improved opportunity for the herd to recover from the invasion.
Refer to captionRefer to caption
Figure 9: $i$ vs $s$ “phase portrait” of a CWD infection with 50 different trials for which no intervention is made (Left) and (Right) for which $\rho_{i}W=0.8$ wolves are introduced when the infected numbers reach 30 (1 20th of the carrying capacity of 600). Here K=30K=30(100km2)-1, $A=2000$ km2, $\mu_{e}=0.2$yr-1.
Refer to captionRefer to caption
Figure 10: (Left) $i$ vs $s$ “phase portrait” of a CWD infection with 50 different trials for which intervention is made with $\rho_{i}W=1.5$ wolves introduced when the infected numbers reach 30 (1 20th of the carrying capacity of 600). (Right): Probability of disease survival as a function of time for the three different control scenarios, $\rho_{i}W=0.0$, $\rho_{i}W=0.8$, $\rho_{i}W=1.5$. Here K=30K=30(100km2)-1, $A=2000$ km2, $\mu_{e}=0.2$yr-1, $\rho_{s}W=0.25\rho_{i}W$.
4.2 Probability of Spread
We can get an analytical understanding of some of these results as follows: Suppose there is one infected cervid with a death rate $\mu(=\mu_{i}+\rho_{i}W)$ introduced into a healthy population of size $s$. What is the probability that prions shed by the one introduced infected individual and its interactions with healthy animals will lead to at least one other cervid becoming infected?
The density of prions resulting from shedding of one infected cervid satisfies the differential equation $\frac{dE}{dt}=\frac{\epsilon A_{0}}{A}-\mu_{e}E$ as long as the infected deer is alive $0<t<t_{d}$, and $\frac{dE}{dt}=-\mu_{e}E$ after its death.
$E(t)=\left\{\begin{array}[]{cc}\beta(1-\exp(-\mu_{e}t)),&0<t<t_{d}\\ \beta(1-\exp(-\mu_{e}t_{d}))\exp(-\mu_{e}(t-t_{d}),&t_{d}<t<\infty\end{array}\right$., (30)
where $t_{d}$ is the time of death of the infected invader, and $\beta=\frac{\epsilon A_{0}}{\mu_{e}A}$. The probability $p_{s}$ of the disease spreading to at least one susceptible individual in a total population of $s$ by time $t$ is given by (ignoring all other births or deaths)
$\frac{dp_{s}}{dt}=\Big{(}\gamma_{e}sE(t)+\gamma_{i}{A_{0}\over A}s(i==1)\Big{)}(1-p_{s}),$ (31)
from which it follows that the probability of spread by time $t$ is given by
$p_{s}(t)=1-\exp\Big{(}-\gamma_{e}s\int_{0}^{t}E(\sigma)d\sigma-\gamma_{i}s{A_{0}\over A}(t_{d}-H(t_{d}-t)(t-t_{d}))\Big{)},$ (32)
where $H$ is the Heaviside function, and the probability of eventual spread is
$p_{s}(\infty)=1-\exp\Big{(}-\gamma_{e}s\int_{0}^{\infty}E(\sigma)d\sigma\ -\gamma_{i}s{A_{0}\over A}t_{d}\Big{)}.$ (33)
A direct calculation using (30) yields
$p_{s}(\infty)=1-\exp\Big{(}-\gamma_{e}s\beta t_{d}-\gamma_{i}s{A_{0}\over A}t_{d}\Big{)}.$ (34)
Now, the probability that the infected animal dies at time $t_{d}$ is given by the probability density function $p_{d}(t_{d})=\mu\exp(-\mu t_{d})$, so the probability of spread, taking into account that the time of death $t_{d}$ is random, is
$p_{s}=\mu\int_{0}^{\infty}\Big{(}1-\exp(-\gamma_{e}s\beta t_{d}-\gamma_{i}{A_{0}\over A}t_{d})\Big{)}\exp(-\mu t_{d})dt_{d}=\frac{1}{\frac{\mu}{\gamma_{e}s\beta+\gamma_{i}s{A_{0}\over A}}+1}.$ (35)
In terms of original parameters,
$\frac{\mu}{\gamma_{e}s\beta+\gamma_{i}s{A_{0}\over A}}=\frac{A\mu_{e}(\mu_{i}+\rho_{i}W)}{sA_{0}(\gamma_{e}\epsilon+\gamma_{i}\mu_{e})}.$ (36)
Notice that $p_{s}$ in (35) is a decreasing function of $\mu$ and an increasing function of $s$. Thus, predation of infected cervids is a beneficial thing, since anything that can increase $\mu$, the death rate of infected cervids, will decrease the probability of spread of the infection. In other words, the presence of wolves in a disease-free population helps to prevent the invasion of the disease by an immigrant infected animal. Further, observe that overpopulation of susceptible cervids increases the probability of the spread of the disease, while predation of the susceptible population, which decreases $s$ through $\rho_{s}W$, helps prevent the spread of the disease. Thus, the presence of predators is beneficial for at least these two reasons.
4.3 Probability of Elimination
Because prions survive in the environment, even if all infected animals are eliminated by some control mechanism, there is the possibility that the disease will reemerge some time later. To calculate this probability, let $p_{s,0}(t)$ be the probability that there are $s$ susceptible deer and no infecteds at time $t$, but there is some contamination of the environment with prions, $E$. It follows that
$\frac{dp_{s,0}}{dt}=-\gamma_{e}sE(t)p_{s,0}.$ (37)
(Here, again, we are ignoring births and deaths of healthy animals). $E(t)$, of course, is the density of prions in the environment, which, as long as there are no infecteds, satisfies $\frac{dE}{dt}=-\mu_{e}E(t)$. Consequently,
$E(t)=E_{0}\exp(-\mu_{e}t),$ (38)
where $E_{0}$ is the initial density of prions at time $t=0$, so that
$\frac{dp_{s,0}}{dt}=-\gamma_{e}sE_{0}\exp(-\mu_{e}t)p_{s,0}.$ (39)
It follows that
$\ln(p_{s,0}(t))={\gamma_{e}\over\mu_{e}}sE_{0}(\exp(-\mu_{e}t)-1),$ (40)
and
$\lim_{t\rightarrow\infty}p_{s,0}=\exp(-{\gamma_{e}\over\mu_{e}}sE_{0})$ (41)
is the probability that there will not be another outbreak. Obviously, the probability that there will not be another outbreak goes to zero as ${\gamma_{e}\over\mu_{e}}sE_{0}$ becomes large. This once again shows the risk associated with overpopulation of susceptible deer (large $s$), and the benefit of having wolves or other harvesting measures to help control the population size of healthy deer, even when there are no infected animals present. This is borne out by a recent statistical study that showed that deer populations that experienced high harvest pressure had lower prevalence of CWD [26].
5 Discussion
Here, we performed a bifurcation analysis of a simple deterministic mathematical model that describes the direct and indirect transmission of CWD in a population of deer subject to predation by predators, such as wolves. We then stochastically simulated the system using the Gillespie Algorithm and applied Poisson-thinning to the prion density, as it varies continuously in time. This model investigation demonstrates that introducing predators, such as wolves, can be an effective way to control diseases such as CWD, for which there are no other known effective control strategies. We also demonstrated that this control strategy works even if CWD is already prevalent in the population, as is the case for many cervid populations in the American West and Upper Midwest. We then computed the probability of elimination, or the probability that CWD will not reemerge through indirect environmental transmission, given a completely susceptible population, and showed that it depends inversely on the size of the susceptible population. The probability that there will not be another CWD outbreak goes to zero exponentially as a function of the susceptible population size. This argues for the need to control not only the size of the infected population, but also the size of the susceptible population. Taken together, this analysis highlights wolves, or other predators, as an effective control strategy for CWD and makes clear the role for human-hunting or other harvesting measures to limit the susceptible population size.
Our analysis has several limitations relating to the parameter values used, model complexity, and modeling framework. We used some published parameter values for our mathematical model, but there are some questions concerning the structural identifiability of some of the model parameters. For a system of this complexity, estimating meaningful parameters from the scarce available data is an ongoing challenge. Consider, for example, a model examined in [24] that has the form
$\text{susceptible Deer Density, Unobserved}~~\dfrac{dS}{dt}$ $=a-S(\beta I+\gamma E+m),$
$\text{Infected Deer Density, Unobserved}~~\dfrac{dI}{dt}$ $=S(\beta I+\gamma E)-I(m+\mu),$
$\text{Prion Density, Unobserved}~~\dfrac{dE}{dt}$ $=\epsilon I-\tau E,$
$\text{Density of Dead Infected Deer, Observed}~~\dfrac{dC}{dt}$ $=\mu I.$
The authors attempted to fit this model using observations of $C$ from two data sets of CWD outbreaks. $S$, II, and $E$, are all unobserved states of the model. A structural identifiability analysis  [^2] We used StructuralIdentifiability.jl to perform the identifiability analysis of the system [9]. The implementation we used can be found here https://docs.sciml.ai/StructuralIdentifiability/stable/. of the model structure reveals that the parameters $\gamma$ and $\epsilon$ are structurally non-identifiable when only $C$ is observed, though the other kinetic parameters are globally identifiable. If $E$ is unobserved, then it should not be expected that $\gamma$ and $\epsilon$ can be uniquely determined based on the structure of the model alone. The parameters $\gamma$ and $\epsilon$ are still globally non-identifiable even if $S$, II, and $C$ are observed, but $E$ is unobserved. However, if one is able to observe $C$ and $E$, but not $S$ and $I$, all kinetic parameters are globally identifiable, but measuring $E$ is a key challenge.
The prions responsible for CWD are known to exist in soil, and can be measured from soil samples [19]. In practice, estimating $E$ through soil samples would likely be expensive, infeasible on a large scale, and miss some of the sources of the prions in the environment. The prions responsible for CWD can be found in multiple environmental reservoirs beyond soil. It was recently shown that ticks can harbor levels of prions that could cause transmission of CWD [17]. Prions were also recently shown to exist in plants at relevant levels for transmission of CWD [7]. For these reasons, getting estimates of the prion density will be difficult. A different approach to ensure that unique parameters can be found is to reparameterize the model to remove the structural identifiability issue ahead of parameter estimation (for an example in a different modeling context, see [11]), but this can alter the interpretability of state variables of the model or rate parameters. Of course, whether unique parameter estimates are needed depends on how the model will be used.
We sought to limit the complexity of the model considered here, largely to keep the analytic calculations for the bifurcation analysis tractable. Given the new findings about the wide range of environmental reservoirs in which prions are found, our model may be overly simple. Both ticks and plants have complex, multi-year seasonal life cycles that exhibit phases of dormancy that are cued by environmental signals, potentially inducing a periodic and stochastic component to indirect environmental infection. These complex dynamics are beyond the scope of our model, but may be important to understanding the spread of CWD in deer at the level of an ecosystem.
We chose to use an ODE model and did not consider the role of possible spatial inhomogeneity. However, the role of space may be important in understanding the spread of CWD in deer. Currently, little is known about the spatial distribution of prions. One would imagine the spatial distribution of the prion is somewhat “patchy” in space because an infected animal carcass contains a high concentration, which would then diffuse in the local environment. However, some degree of spatial mixing is likely to occur. For example, rivers could be an agent by which the prions could spread in the environment. Prions can remain infectious after passing through a crow digestive system [10, 38] and a coyote digestive system [27]. The same may be expected from a wolf, and so it may be expected that wolves would spread the prions throughout the environment through feces. In our model, we took the term $\xi W$ in Eq. 9 to be zero throughout our analysis, but taking this parameter to be non-zero would begin to capture prion spread by wolves. Currently, less is known about the spread of prions through insects and plants, or if infected plants can pass the prions inter-generationally through seeds, but these are other theoretical mechanisms by which prions could become well-mixed in the environment. More research is needed to understand the spatial distribution of prions in the environment and how this contributes to the indirect transmission of CWD.
This modeling analysis shows that predators can be used as an effective control strategy for managing the spread of CWD in deer populations and highlights a possible role for hunting or other harvesting of the susceptible population to increase the probability of CWD elimination. However, this analysis should be understood as a proof-of-concept and not used for management purposes, given the uncertainty that exists in the parameters of the system. Inferring biologically relevant parameters, such as the indirect transmission rate and the growth rate of $E$, from the available data is a key challenge for this system. As more is understood about the dynamics of environmental reservoirs of prions, such as ticks and plants, the spatial distribution of prions in the environment, and, more broadly, an ecosystem-scale understanding of CWD transmission, our model can be modified to understand these dynamical processes.
6 Funding
C. E. F. is supported in part by The James S. McDonnell Foundation Postdoctoral Fellowship Award in Complex Systems (https://doi.org/10.37717/2020-1591) and by the NSF-Simons Center for Quantitative Biology at Northwestern University (NSF: 1764421 and Simons Foundation/SFARI 597491-RWC).
7 Author Contribution
C.E.F. and J.P.K. designed the research. J.P.K. performed the model analysis and simulations. C.E.F. reviewed the analysis and performed the identifiability analysis. J.P.K. and C.E.F. wrote and edited the paper.
8 Code
The code used for this analysis can be found here https://github.com/cefitzg/cwd_code.git